OpenStudy (therealmeeeee):

I need a little help here Factor. x2 – 6x + 9 A. (x – 3)2 B. (x + 3)2 C. (x – 9)2 D. (x + 9)2

OpenStudy (anonymous):

x^2-6x+9 = (x-3)(x-3) = (x-3)^2 12x^2+x-6 a=12 b=1 c=-6 and ac=-72 12x^2+x-6 = 12x^2 +9x-8x-6= 3x(4x+3)-2(4x+3) = (4x+3)(3x-2)

OpenStudy (anonymous):

A